OpenStudy (anonymous):

Ten grams of silver will require calories to raise the temperature 1° C. silver has 0.06 specific heat

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

ten rams of silver will require how many calories

OpenStudy (anonymous):

do you know the formula? is this a physics or chemistry question?

OpenStudy (anonymous):

chemistry and its called specific heats

OpenStudy (anonymous):

\[mc \Delta t\]

OpenStudy (anonymous):

q=mc(delta)t

OpenStudy (anonymous):

so does your teacher want you to convert to SI units?

OpenStudy (anonymous):

because we have to convert temperature to kelvin and mass to kilogram

OpenStudy (anonymous):

mass=10/1000=0.01 kg temperature=1+273 = 274Kelvin specific heat=0.06j/kgk Q=0.01*0.06*274 =0.1644Joules

OpenStudy (anonymous):

the quantity of heat that ten grams of silver requires

OpenStudy (anonymous):

and you can easily convert from joules to calories...